A company that manufactures memory chips for digital cameras uses the formula c = 129 ^3√n^2 + 1295 to determine the cost c, in dollars, of producing n chips. How much will it cost to produce 226 chips?

Answer by stanbon(75887)

A company that manufactures memory chips for digital cameras uses the formula c = 129 ^3√n^2 + 1295 to determine the cost c, in dollars, of producing n chips. How much will it cost to produce 226 chips?
C(226) = 129^3*sqrt(226^2)+1295 = 485153009
----
Comment:: I doubt it.
Cheers,
Stan H.